# On Utility Maximization under Multivariate Fake Stationary Affine Volterra Models

### 06 Apr 2026 v2

@@ 2. $(b)$
where the last term is finite due to condition (3.27) and the inequality $|z|^{q}\leq c_{q}e^{|z|},\;\forall\,q\geq 1$. It becomes straightforward that $\alpha^{*}$ in equation (3.78) is admissible.
Step 2 (Proof of the equality (1):) Let’s us consider the Problem (3.57) with an arbitrary strategy $\alpha\in\mathcal{A}$. To ease notations, we set $Y_{t}:=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}\right)$ and $h_{t}=\lambda_{t}+\Sigma\Lambda_{t}$. For any admissible strategy $\alpha\in\mathcal{A}$, Itô’s lemma combined with the property of $\Gamma$ in Equation (3.72) of Proposition 3.11 yield:
$d\big(\Gamma_{t}Y_{t}\big)\;$ $=\Gamma_{t}Y_{t}\Big(D_{t}(\alpha_{t})dt-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}+\Lambda_{t}^{\top}dW_{t}\Big).$
where the drift factor takes the form $D_{t}(\alpha_{t})=\frac{\gamma^{2}}{2}e^{\int_{t}^{T}2r(s)ds}\alpha_{t}^{\top}\alpha_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}h_{t}+\frac{1}{2}h_{t}^{\top}h_{t}.$ As a consequence, using $\Gamma_{T}=1$, we get
$\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)e^{\int^{T}_{0}D_{s}(\alpha_{s})ds}F_{T}^{\alpha}.$ (5.103)
where $F_{T}^{\alpha}$ is an $(\mathbb{F},\mathbb{P})$-local martingale as $(\alpha,\Lambda)$ are in $L^{2,{loc}}_{\mathbb{F}}([0,T])$ and write
$F_{T}^{\alpha}$ $=\mathcal{E}\Big(\int_{0}^{T}-\gamma e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)$
$=\operatorname{exp}\Big(-\frac{1}{2}\int^{T}_{0}\Big(\gamma^{2}e^{2\int_{s}^{T}r(u)du}\left|\alpha_{s}\right|^{2}+\left|\Lambda_{s}\right|^{2}\Big)ds-\gamma\int_{0}^{T}e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\int_{0}^{T}\Lambda_{s}^{\top}dW_{s}\Big).$
Now, for $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a $\mathbb{P}$-martingale with expectation $1$ by Lemma 5.1 with $g_{1,i}(t)=-\rho_{i}\theta_{i}+(1-\rho_{i}^{2})\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\in L^{\infty}([0,T],\mathbb{R})$ and $g_{2,i}(t)=-\sqrt{1-\rho_{i}^{2}}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\in L^{\infty}([0,T],\mathbb{R})$ for $1\leq i\leq d$. Consequently, inserting the candidate (3.78) for the optimal strategy $\alpha^{*}$ in the above Equation (5.103), and observing that $D_{s}(\alpha_{s}^{*})=0\;\forall\,s\in[0,T]$, then taking the expectation, its boils down that:
$\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha^{*}}_{T}\right)\big]}=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}{\big[F_{T}^{\alpha^{*}}\big]}=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right).$
where the last inequality comes from the fact that $F_{T}^{\alpha^{*}}$ is a true martingale. This completes the first part $(1)$ of the proof. It remains to show the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$.
Step 3 (Proof of the inequality (2):) Note that the SDE for the wealth process (3.55) can be solved explicitly so as to admit the representation
$X_{T}^{\alpha}=e^{\int^{T}_{0}r(s)ds}x_{0}+\int^{T}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}\mathrm{d}s+\int^{T}_{0}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}dB_{s}.$ (5.104)
Since $\alpha_{t}$ is bounded by assumption and $s\to e^{\int^{T}_{s}r(u)du}\in L^{\infty}([0,T],\mathbb{R})$, we can define a new probability measure $\mathbb{Q}$ with Radon-Nikodym density
${Z}_{T}:=\frac{d{\mathbb{Q}}}{d\mathbb{P}}|_{\mathcal{F}_{T}}=\exp\Big(-\gamma\int_{0}^{T}e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\,dB_{s}-\frac{\gamma^{2}}{2}\int_{0}^{T}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}ds\Big),$ (5.105)
by Lemma 5.2. Recalling that $U(X_{T}^{\alpha})=-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)$, by Girsanov’s theorem, we may write:
$\mathbb{E}_{x_{0},V_{0}}\Big[U(X_{T}^{\alpha})\Big]$ $\,=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}\big(-\gamma e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}\big)ds\Big)\Big]$
$=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}^{\alpha}ds\Big)\Big]$
with the real-valued deterministic process $F_{s}$ given by
$\forall\,s\in[0,T],\quad F_{s}^{\alpha}$ $=-\gamma e^{\int^{T}_{s}r(u)du}\alpha_{s}^{\top}\lambda_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\alpha_{s}\right|^{2}$
$=F_{s}^{\alpha^{*}}-\gamma e^{\int^{T}_{s}r(u)du}\hat{\alpha}_{s}^{\top}\lambda_{s}+\gamma^{2}e^{2\int^{T}_{s}r(u)du}\alpha_{s}^{*\top}\hat{\alpha}_{s}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\left|\hat{\alpha}_{s}\right|^{2}$
where we write the arbitrary admissible strategy $(\alpha_{t})_{t\in[0,T]}$ in terms of the optimal strategy $\alpha^{*}_{t}=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)$ for every $t\in[0,T]$ and some remainder $(\hat{\alpha}_{t})_{t\in[0,T]}$ i.e. $\alpha_{t}=\alpha_{t}^{*}+\hat{\alpha}_{t}$ for every $t\in[0,T]$. Note that, on the first hand, substituting the optimal strategy $\alpha^{*}$ in (3.78) yields $\forall\,s\in[0,T]$
$F_{s}^{\alpha^{*}}=-\left(\lambda_{s}+\Sigma\Lambda_{s}\right)^{\top}\lambda_{s}+\frac{1}{2}\left|\lambda_{s}+\Sigma\Lambda_{s}\right|^{2}=\frac{1}{2}(-\left|\lambda_{s}\right|^{2}+\left|\Sigma\Lambda_{s}\right|^{2})=\sum_{i=1}^{d}\Big(-\frac{\theta_{i}^{2}}{2}+\frac{1}{2}\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}$
and on the other hand, recalling that $\alpha_{s}:=\sqrt{\operatorname{diag}(V_{s})}^{\top}\pi_{s}$ and defining $\hat{\pi}:=\pi-\pi^{*}$ so that $\pi_{s}=\pi_{s}^{*}+\hat{\pi}_{s}$, with $\pi_{s}^{*}=\frac{1}{\gamma}e^{-\int_{s}^{T}r(u)du}\left(\theta+\Sigma\nu\varsigma(s)\psi(T-s)\right)$ so that $\pi_{s,i}^{*}=\frac{1}{\gamma}e^{-\int_{s}^{T}r(u)du}(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s))$ for $i=1,\ldots,d$ and consequently, $\forall\,s\in[0,T],$
$F_{s}^{\alpha}$ $=F_{s}^{\alpha^{*}}+\sum_{i=1}^{d}\Big(-\gamma e^{\int^{T}_{s}r(u)du}\theta_{i}\hat{\pi}_{s,i}+\gamma^{2}e^{2\int^{T}_{s}r(u)du}\pi_{s,i}^{*}\hat{\pi}_{s,i}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
$=\sum_{i=1}^{d}\Big(-\frac{\theta_{i}^{2}}{2}+\frac{1}{2}\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}+\gamma e^{\int^{T}_{s}r(u)du}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
Using Equation (3.67) reading $\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))=(\psi^{i}\star r^{i})^{\prime}(T-s)$ owing to Lemma 5.3 $(a)$ for $i=1,\ldots,d$, where $r^{i}$ the resolvent of the first kind of the integral kernel $K_{i}$, one gets:
$\forall\,s\in[0,T],\;F_{s}$ $=\sum_{i=1}^{d}(\psi^{i}\star r^{i})^{\prime}(T-s)V^{i}_{s}-\Big(F_{i}(s,\psi(T-s))-\frac{1}{2}\rho_{i}^{2}\nu_{i}^{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V^{i}_{s}$ (5.106)
$\hskip 42.67912pt+\sum_{i=1}^{d}\Big(\gamma e^{\int^{T}_{s}r(u)du}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}+\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}\Big)V^{i}_{s}$
Under the probability measure $\mathbb{Q}$ defined in (5.105), the processes $\widehat{B}_{t}:=B_{t}+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\alpha_{s}ds$ and
$\widehat{W}_{t}:=\Sigma\widehat{B}_{t}+\sqrt{I-\Sigma^{\top}\Sigma}B^{\perp}_{t}=W_{t}+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\Sigma\alpha_{s}ds=W_{t}+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\Sigma\alpha^{*}_{s}ds+\gamma\int_{0}^{t}e^{\int^{T}_{s}r(u)du}\Sigma\hat{\alpha}_{s}ds,$
are standard Wiener processes thanks to the Cameron–Martin–Girsanov and Girsanov theorem respectively. The dynamics of the variance process $V$ under $\mathbb{Q}$ can thus be written as
$H_{s}^{i}:=(DV_{s})_{i}-\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)e^{\int^{T}_{s}r(u)du}(\pi_{s,i}^{*}+\hat{\pi}_{s,i})V_{s}^{i},\quad i=1,\ldots,d.$
$\hskip 17.07182pt=-\rho_{i}\theta_{i}\nu_{i}\varsigma^{i}(s)V_{s}^{i}+(DV_{s})_{i}-(\rho_{i}\nu_{i}\varsigma^{i}(s))^{2}\psi^{i}(T-s)V_{s}^{i}-\gamma\rho_{i}\nu_{i}\varsigma^{i}(s)e^{\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}V_{s}^{i}$
$\hskip 42.67912pt=\sum_{i=1}^{d}\int_{t}^{T}\big(-\frac{1}{2}\theta_{i}^{2}+F_{i}(s,\psi(T-s))\big)g_{0}^{i}(s)ds+\sum_{i=1}^{d}\int_{t}^{T}\psi^{i}(T-s)\big(H_{s}^{i}ds+\nu_{i}\varsigma^{i}(s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\big).$
Recalling the change of variable in (5.98), one gets using Equation (3.70)
$\sum_{i=1}^{d}\int_{0}^{T}\psi^{i}(T-s)H_{s}^{i}ds$ $=\sum_{i=1}^{d}\int_{0}^{T}\Big(F_{i}(s,\psi^{i}(T-s))-\frac{\nu_{i}^{2}}{2}\big(1+\rho_{i}^{2}\big)(\varsigma^{i}(s)\psi^{i}(T-s))^{2}\Big)V_{s}^{i}\,ds$
$-\gamma\sum_{i=1}^{d}\int_{0}^{T}e^{\int^{T}_{s}r(u)du}\rho_{i}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\hat{\pi}_{s,i}V_{s}^{i}ds.$
Replacing back all in (5.106), some terms cancel out and we end up with
$\int_{0}^{T}F_{s}\,ds$ $=\sum_{i=1}^{d}\int_{0}^{T}\Big(-\frac{1}{2}\theta_{i}^{2}+F_{i}(s,\psi(T-s))\Big)g_{0}^{i}(s)\,ds+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds$
$\qquad+\sum_{i=1}^{d}\int_{0}^{T}\Big(-\frac{\nu_{i}^{2}}{2}\big(\varsigma^{i}(s)\psi^{i}(T-s)\big)^{2}V_{s}^{i}\,ds+\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}\,d\widehat{W}^{i}_{s}\Big).$
Hence we obtain
$\,\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)\Big]=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(\int_{0}^{T}F_{s}^{\alpha}ds\Big)\Big],$
$\hskip 128.0374pt=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 79.6678pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}$
$\hskip 170.71652pt+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds\Big)\Big].$ (5.107)
$\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha^{*}}_{T}\right)\Big]=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 62.59596pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}^{*}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{*i}_{s}\Big)\Big].$
Since $\psi$ is continuous, it is bounded and therefore the stochastic exponential is a true $\mathbb{Q}^{*}$- martingale with expectation $1$ by Lemma 5.1 with $g_{2}=0$ and $g_{1,i}(s)=\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)$ . Thus we get
$\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha^{*}}_{T}\right)\Big]=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$
This completes the first part of the proof (equality (1)).
Back to the proof of the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$, still from Equation (5.107), since $V_{s}$ is positive definite, the term $\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds$ or equivalently $\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma^{2}}{2}e^{2\int^{T}_{s}r(u)du}\hat{\alpha}_{s,i}^{2}\,ds$ has to be more than or equal to 0 so that
$\mathbb{E}_{x_{0},V_{0}}\Big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)\Big]\leq-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big)$
$\hskip 62.59596pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\Big)\Big].$
$\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)\big]}\leq-\frac{1}{\gamma}\exp\left(-\gamma e^{\int^{T}_{0}r(s)ds}x_{0}\right)\exp\Big(\sum_{i=1}^{d}\int_{0}^{T}\big(-\frac{\theta_{i}^{2}}{2}+F_{i}(s,\psi(T-s))\big)g^{i}_{0}(s)ds\Big).$
which completes the proof and we are done. $\Box$
@@
In these terms we are bound to choose a function $f$ for which $J_{t}^{\alpha}$ is a supermartingale for all $\alpha\in\mathcal{A}$ and there exists a $\alpha^{*}\in\mathcal{A}$ such that $J_{t}^{\alpha^{*}}$ is a martingale. We assume that $(\Gamma,\Lambda)\in\mathbb{S}^{\infty}_{\mathbb{F}}([0,T],\mathbb{R}^{*}_{+})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$. Note that $J_{t}^{\alpha}$ satisfies conditions $1$ and $2$ of Definition 3.2 since $\Gamma_{T}=0$ and $\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$. Applying Itô’s formula on $J_{t}^{\alpha}$ yields $dJ_{t}^{\alpha}=\big(r(t)+\alpha_{t}^{\top}\lambda_{t}-\tfrac{1}{2}\left|\alpha_{t}\right|^{2}-f(t,\Gamma_{t},\Lambda_{t})\big)\,dt+\alpha_{t}^{\top}\,dB_{t}+\Lambda_{t}^{\top}dW_{t}$. In order for the third condition to be fulfilled by $J_{t}^{\alpha}$, we may have that $r(t)+\alpha_{t}^{\top}\lambda_{t}-\tfrac{1}{2}\left|\alpha_{t}\right|^{2}\leq f(t,\Gamma_{t},\Lambda_{t}),$ for all $t\in[0,T]$ and $\alpha\in\mathcal{A}$. The pointwise maximization yields $\alpha^{*}_{t}:=\lambda_{t}$ for every $t\in[0,T]$ and $dM_{t}:=\lambda_{t}^{\top}\,dB_{t}+\Lambda_{t}^{\top}dW_{t}$ is a true martingale by condition (3.27)( since Assumption 3.1 holds). The admissibility of $\alpha^{*}$ is straightforward thanks to the same assumptions. The corresponding value function is determined simply by the initial value of the supermartingale $J_{t}^{\alpha}$ at $\alpha^{*}$. Subtituting $\alpha^{*}$ into (5.108) and taking the expectation yields:
@@
$\hskip 142.26378pt\forall\,t\in[0,T]\quad J_{t}^{\alpha}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}e^{\int_{0}^{t}D_{s}(\alpha_{s})\,ds}F_{t}^{\alpha}\quad\text{where}$ (A.111)
$F_{t}^{\alpha}=\mathcal{E}\big(\int_{0}^{t}\gamma\,\alpha_{s}^{\top}\,dB_{s}+\delta\,\Lambda_{s}^{\top}\,dW_{s}\big)=\operatorname{exp}\Big(-\frac{1}{2}\int^{t}_{0}\big(\gamma^{2}\left|\alpha_{s}\right|^{2}+\delta^{2}\left|\Lambda_{s}\right|^{2}\big)ds+\gamma\int_{0}^{t}\alpha_{s}^{\top}dB_{s}+\delta\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big).$
Now, since $D_{s}(\alpha_{s})\leq 0$, $e^{\int^{t}_{0}D_{s}(\alpha_{s})ds}$ is a non-increasing function. By our assumptions on the admissible strategies 3.3 and Proposition 3.4, $\left(\alpha,\Lambda\right)\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ and thus the stochastic exponential $F_{t}^{\alpha}$ is a local martingale (which follows from the basic properties of the Dooléans-Dade exponential). Therefore, there exists a sequence of stopping times $\{\tau_{n}\}_{n\geq 1}$ satisfying $\lim_{n\rightarrow\infty}\tau_{n}=T$, $\mathbb{P}$-a.s., such that
@@
For $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a martingale by Lemma 5.1 with $g_{1,i}(t)=\frac{\gamma}{1-\gamma}\rho_{i}\theta_{i}+(1+\frac{\gamma}{1-\gamma}\rho_{i}^{2})\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\in L^{\infty}([0,T],\mathbb{R})$ and $g_{2,i}(t)=\frac{\gamma}{1-\gamma}\sqrt{1-\rho_{i}^{2}}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\in L^{\infty}([0,T],\mathbb{R})$ for $1\leq i\leq d$. Subsequently, $J^{\alpha^{*}}_{t}$ is a true martingale. We have verified all conditions required by martingale optimality principle, except for the admissibility of $\alpha^{*}$, which follows directly from Step 1 in the proof of Theorem 3.7 in the more general setting, as the arguments are similar. The proof is complete. $\Box$
@@
$\text{where}\quad F_{t}^{\alpha}=\mathcal{E}\Big(\int_{0}^{t}-\gamma e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)$
$\hskip 56.9055pt=\operatorname{exp}\Big(-\frac{1}{2}\int^{t}_{0}\Big(\gamma^{2}e^{2\int_{s}^{T}r(u)du}\left|\alpha_{s}\right|^{2}+\left|\Lambda_{s}\right|^{2}\Big)ds-\gamma\int_{0}^{t}e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big).$
Note $D_{t}(\alpha^{*}_{t})=0$, then $D_{t}(\alpha_{t})\geq 0$. By our assumptions on the admissible strategies 3.8 and Proposition 3.9, $\left(\alpha,\Lambda\right)\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ and thus, the stochastic exponential $F_{t}^{\alpha}$ is a local martingale (which follows from the basic properties of the Dooléans-Dade exponential). Therefore, there exists a sequence of stopping times $\{\tau_{n}\}_{n\geq 1}$ satisfying $\lim_{n\rightarrow\infty}\tau_{n}=T$, $\mathbb{P}$-a.s., such that that $F^{\alpha}_{t\wedge\tau_{n}}$ is a positive martingale for every $n$.

### 07 May 2026 v3

@@ 2. $(b)$
where the last term is finite due to condition (3.29) and the inequality $|z|^{q}\leq c_{q}e^{|z|},\;\forall\,q\geq 1$. It becomes straightforward that $\alpha^{*}$ in equation (3.81) is admissible. This complete the Proof $\Box$
@@
In these terms we are bound to choose a function $f$ for which $J_{t}^{\alpha}$ is a supermartingale for all $\alpha\in\mathcal{A}$ and there exists a $\alpha^{*}\in\mathcal{A}$ such that $J_{t}^{\alpha^{*}}$ is a martingale. Note that $J_{t}^{\alpha}$ satisfies conditions $1$ and $2$ of Definition 3.2 since $\Gamma_{T}=0$ and $\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$. Applying Itô’s formula on $J_{t}^{\alpha}$ yields
$dJ_{t}^{\alpha}=\big(r(t)+\alpha_{t}^{\top}\lambda_{t}-\tfrac{1}{2}\left|\alpha_{t}\right|^{2}-f(t,\Gamma_{t},\Lambda_{t})\big)\,dt+\alpha_{t}^{\top}\,dB_{t}+\Lambda_{t}^{\top}dW_{t}$
. In order for the third condition to be fulfilled by $J_{t}^{\alpha}$, we may have that
$r(t)+\alpha_{t}^{\top}\lambda_{t}-\tfrac{1}{2}\left|\alpha_{t}\right|^{2}\leq f(t,\Gamma_{t},\Lambda_{t}),\quad\text{for all}\;t\in[0,T]\;\text{and}\;\alpha\in\mathcal{A}.$
The pointwise maximization yields $\alpha^{*}_{t}:=\lambda_{t}$ for every $t\in[0,T]$. We define the driver of the BSDE (5.110) by $f(t):=r(t)+\frac{1}{2}|\lambda_{t}|^{2}$ for $t\in[0,T]$, which is $\mathcal{F}_{t}$-measurable. Classical results on BSDEs (see for example Zhang (2017)) yield the existence and uniqueness of a solution $(Y,\Lambda)\in\mathbb{S}^{2}_{\mathbb{F}}([0,T],\mathbb{R})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$ to (5.110) so that $dM_{t}:=\lambda_{t}^{\top}\,dB_{t}+\Lambda_{t}^{\top}dW_{t}$ is a true martingale by Lemma 5.1 ( since Assumption 3.1 is in force). The admissibility of $\alpha^{*}$ is straightforward thanks to the same assumptions (see condition (3.29)). The corresponding value function is determined simply by the initial value of the supermartingale $J_{t}^{\alpha}$ at $\alpha^{*}$. Subtituting $\alpha^{*}$ into (5.109) and taking the expectation yields:
@@
$\hskip 28.45274pt\forall\,t\in[0,T]\quad J_{t}^{\alpha}=\frac{\Gamma_{0}x_{0}^{\gamma}}{\gamma}e^{\int_{0}^{t}D_{s}(\alpha_{s})\,ds}F_{t}^{\alpha}$ (A.112)
$\quad\text{where}\quad F_{t}^{\alpha}=\mathcal{E}\big(\int_{0}^{t}\gamma\,\alpha_{s}^{\top}\,dB_{s}+\delta\,\Lambda_{s}^{\top}\,dW_{s}\big).$ (A.113)
Now, since $D_{s}(\alpha_{s})\leq 0$, $e^{\int^{t}_{0}D_{s}(\alpha_{s})ds}$ is a non-increasing function. By our assumptions on the admissible strategies 3.3 and Proposition 3.4, $\left(\alpha,\Lambda\right)\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})^{2}}$ and thus the stochastic exponential $F_{t}^{\alpha}$ is a local martingale (which follows from the basic properties of the Dooléans-Dade exponential). Therefore, there exists a sequence of stopping times $\{\tau_{n}\}_{n\geq 1}$ satisfying $\lim_{n\rightarrow\infty}\tau_{n}=T$, $\mathbb{P}$-a.s., such that
@@
For $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a martingale by Lemma 5.1 with $g_{1}(t)=\frac{\gamma}{1-\gamma}$ , $g_{2}(t)=\delta$ for every $t\in[0,T]$ and $\kappa=\delta$, Subsequently, $J^{\alpha^{*}}_{t}$ is a true martingale. We have verified all conditions required by martingale optimality principle, except for the admissibility of $\alpha^{*}$, which follows directly from Step 1 in the proof of Theorem 3.7 in the more general setting, as the arguments are similar. The proof is complete. $\Box$
@@
$\text{where}\quad F_{t}^{\alpha}=\mathcal{E}\Big(\int_{0}^{t}-\gamma e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big).$
Note $D_{t}(\alpha^{*}_{t})=0$, then $D_{t}(\alpha_{t})\geq 0$. By our assumptions on the admissible strategies 3.8 and Proposition 3.9, $\left(\alpha,\Lambda\right)\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})^{2}}$ and thus, the stochastic exponential $F_{t}^{\alpha}$ is a local martingale (which follows from the basic properties of the Dooléans-Dade exponential). Therefore, there exists a sequence of stopping times $\{\tau_{n}\}_{n\geq 1}$ satisfying $\lim_{n\rightarrow\infty}\tau_{n}=T$, $\mathbb{P}$-a.s., such that that $F^{\alpha}_{t\wedge\tau_{n}}$ is a positive martingale for every $n$.